OpenStudy (anonymous):

how do i simplify -3x - 9 + 15x

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

once you've combined like terms you can factor which should simplify as much as you can.

OpenStudy (anonymous):

thank you

OpenStudy (shoba):

-3x - 9 + 15x = 12x - 9

OpenStudy (anonymous):

you dont have to factor just combine like terms

OpenStudy (anonymous):

thnank you shoba

OpenStudy (anonymous):

they're asking to simplify not factor if the equation had x^2 then yes it would possibly be factor

OpenStudy (anonymous):

Right @shoba

OpenStudy (anonymous):

another question , -9k + 8k?

OpenStudy (anonymous):

add 8 to -9

OpenStudy (shoba):

-3x - 9 + 15x = 12x - 9 = 3(4x -3) -9k + 8k = -k